OpenStudy (anonymous):

can some one help me solve the equation f(x)=5x+12. find f-1(x)

OpenStudy (zzr0ck3r):

\(y=5x+12\) one approach is to switch the \(x'a \text{ and } y's\) and then solve for \(y\). \(y=5x+12 \rightarrow x=5y+12 \rightarrow x-12=5y \rightarrow y = \frac{x-12}{5}\) so \[f^{-1}(x)=\frac{x-12}{5}\]

OpenStudy (zzr0ck3r):

that should say "one approach is to switch the x's and y's"

OpenStudy (anonymous):

thnx